# CHASE: A Causal Heterogeneous Graph based Framework for Root Cause Analysis in Multimodal Microservice Systems

### v1

@@ V Conclusion
For complex microservice systems, conducting a root cause analysis is essential in identifying and addressing the underlying issues that hinder the optimal performance of the system. To tackle this challenge, we propose a causal heterogeneous graph based framework named CHASE to accurately locate the root cause instances of failures, facilitating effective troubleshooting and system maintenance. Starting with modeling microservice systems with graph representations, we apply encoders and embedding layers combined with positional encoding to encode the multimodal data of trace, log and metrics into an invocation graph. CHASE leverages heterogeneous message passing on invocation graphs to tackle instance-level anomaly detection. Next, it constructs a hypergraph with each hyperedge capturing the causality flow for each instance. Finally, the designated hypergraph module locates the root cause anomaly with a detection head applied to each updated node embedding after hypergraph convolution. We comprehensively evaluate CHASE using two real-world datasets, and the evaluation results confirm the superiority of our proposed framework. For the future work, hypergraph learning techniques on modelling spatiotemporal graphs can be applied to capture the causality structure of the trace, which is dynamically evolving since the whole trace typology is gradually formed by sequential invocations ranging in a certain time interval. Additionally, a single powerful autoregressive encoder that is intrinsically trained on multimodal data, such as the large language model, can substitute the dedicated log and metric encoders of CHASE, in order to enhance the quality of embeddings, as well as robustly unify the information of metrics and log.

### v2

@@ V Conclusion
For complex microservice systems, conducting a root cause analysis is essential in identifying and addressing the underlying issues that hinder the optimal performance of the system. To tackle this challenge, we propose a causal hypergraph based framework named CHASE to accurately locate the root cause instances of failures, facilitating effective troubleshooting and system maintenance. Starting with modeling microservice systems with graph representations, we apply encoders and embedding layers combined with positional encoding to encode the multimodal data of trace, log and metrics into an invocation graph. CHASE employs heterogeneous message passing on invocation graphs to tackle instance-level anomaly detection. Next, it constructs a hypergraph with each hyperedge capturing the causality flow for each instance. Finally, the designated hypergraph module locates the root cause anomaly with a detection head applied to each updated node embedding after hypergraph convolution. We evaluate CHASE using two real-world datasets with seven baseline methods, and the evaluation results confirm the superiority of our proposed framework. For the future work, hypergraph learning techniques on modelling spatiotemporal graphs can be applied to capture the causality structure of the trace, which is dynamically evolving since the whole trace typology is gradually formed by sequential invocations ranging in a certain time interval. Additionally, a single powerful autoregressive encoder that is intrinsically trained on multimodal data, such as the large language model, can substitute the dedicated log and metric encoders of CHASE, in order to enhance the quality of embeddings, as well as robustly unify the information of metrics and log.